OpenStudy (anonymous):

A bike shop sells rolling strollers for runners with babies, with a total of 28 in the store. Some of the strollers have 3 wheels and some have 4. If there are 100 wheels, how many 3-wheel strollers are there? A. 25 B. 12 C. 3 D. 16

OpenStudy (anonymous):

x= 3 wheels y = 4 wheels x+y=28 y=28-x 3x+4y=100 3x+4(28-x)=100 3x+112-4x=100 -x=-112+100 Add 100 to -112 to get -12. -x=-12 multiply by -1 x=12